# Complexity Bounds for Smooth Convex Multiobjective Optimization

### v1

@@ 3.4 Oracle Model and Algorithm Classes
The oblivious one-step class models standard gradient descent with a fixed step schedule. It does not include methods with momentum, such as Nesterov’s accelerated gradient method (AGD) or Polyak’s heavy-ball method.
@@
Moreover, if the method is an oblivious one-step gradient method on $g$, i.e.,
@@
Inductive step. Assume (4) holds for all indices up to $t$. By the span property (2), there exist (possibly data-dependent) scalars $\{\beta_{j}^{(t)}\}_{j=0}^{t}$ such that
@@
Subtracting $x^{\star}$ and using (3) gives
@@ 4 Oracle Complexity Lower Bounds
We present our lower bounds by constructing "resisting oracle" instances. In all constructions, we define the initial distance to the Pareto set as $R:=\mathrm{dist}(x^{(0)},\mathcal{P})$. For any specific scalarization $\lambda\in\Delta^{m}$, the distance to its unique minimizer $x_{\lambda}^{\star}$ satisfies $R_{\lambda}:=\|x^{(0)}-x^{\star}_{\lambda}\|\geq R$. Our hard instances are constructed such that we can choose a $\lambda$ for which $R_{\lambda}=R$, making the bounds tight.
4.1 The Strongly Convex Case: A Tight Linear Rate
We begin by establishing a lower bound on the convergence rate for strongly convex multiobjective optimization. Our analysis considers the broad class of span first-order methods and demonstrates that any such algorithm is subject to a linear convergence rate fundamentally limited by the problem’s condition number. The following theorem formalizes this result.
Theorem 4.1 (Strongly Convex Lower Bound).
Fix $L\geq\mu>0$, $T\in\mathbb{N}$, $\kappa:=L/\mu$, and $R>0$. Consider any span first-order method that makes $T$ first-order oracle calls when applied to a quadratic objective. Then there exists a symmetric positive definite matrix $H\in\mathbb{R}^{d\times d}$ with spectrum $\sigma(H)\subset[\mu,L]$ (for some $d\geq T+1$), a vector $b\in\mathbb{R}^{d}$, and an initialization $x^{(0)}\in\mathbb{R}^{d}$ such that for the quadratic
$g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x,\qquad x^{\star}=H^{-1}b,\qquad\|x^{(0)}-x^{\star}\|=R,$
the $T$-th iterate produced by the method satisfies
$\bigl{\|}\nabla g\bigl{(}x^{(T)}\bigr{)}\bigr{\|}\;\geq\;\mu\,R\;\frac{2}{\rho^{T}+\rho^{-T}}\;\geq\;\mu\,R\left(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\right)^{T},\qquad\rho:=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$
In particular, any such method has worst-case iteration complexity $\Omega\!\bigl{(}\sqrt{\kappa}\log(1/\varepsilon)\bigr{)}$ to reduce the gradient norm below $\varepsilon$.
@@ 4.1 The Strongly Convex Case: A Tight Linear Rate
Let $g(x)=\frac{1}{2}x^{\top}Hx-b^{\top}x$ with $H\succ 0$. For any iterate sequence $\{x^{(t)}\}$ produced by a span first-order method on a quadratic, the error admits the polynomial representation (Lemma 3.13):
@@
where the last inequality follows from $\frac{2}{a+a^{-1}}\geq a^{-1}$ for $a\geq 1$. The iteration-complexity statement follows by solving $\mu R\bigl{(}(\sqrt{\kappa}-1)/(\sqrt{\kappa}+1)\bigr{)}^{T}\leq\varepsilon$ for $T$. ∎
Remark 4.2 (Lifting to MOO).
The construction of Theorem 4.5 with $\gamma=\mu$ yields $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ and $\mathrm{dist}(x^{(0)},\mathcal{P})=R$, thus Theorem 4.1 extends directly to MOO.
4.2 The Convex Case: Oblivious vs. Adaptive Methods
We now turn to the smooth convex setting, where strong convexity is no longer assumed. In this regime, the achievable complexity bounds depend crucially on the information available to the algorithm. We therefore draw a distinction between oblivious one-step methods, which employ a fixed scalarization and pre-scheduled step sizes, and fully adaptive span methods. We derive distinct lower bounds for each class, revealing a fundamental separation in their worst-case performance guarantees.

### v2

@@ 3.4 Oracle Model and Algorithm Classes
Definition 3.11 (Oblivious Span First-Order Methods).
A span first-order method is oblivious if, for each $t\geq 1$, there exist coefficients $\{\beta^{(t)}_{i,j}\}_{i=1,\ldots,m;\,j=0,\ldots,t-1}$ that are fixed in advance (possibly as a function of known problem parameters such as $L,\mu,m,T$) and do not depend on the realized oracle replies, such that
$x^{(t)}\;=\;x^{(0)}+\sum_{j=0}^{t-1}\sum_{i=1}^{m}\beta^{(t)}_{i,j}\,\nabla f_{i}\bigl(x^{(j)}\bigr).$
Equivalently, the method satisfies the span property and uses only pre-scheduled (instance-independent) linear combination coefficients.
Remark 3.12.
The class of oblivious span methods includes accelerated schemes with fixed parameter schedules such as Nesterov’s accelerated gradient method on a fixed scalarization.
@@
The oblivious one-step class models standard gradient descent with a fixed step schedule. It does not include methods with momentum, such as AGD or Polyak’s heavy-ball method.
@@
Moreover, if the method is an oblivious span method (Definition 3.11) when applied to $g$, then for each $t$ the polynomial $p_{t}$ depends only on the method’s fixed coefficients (and on $t$), and is therefore independent of the specific quadratic instance $(H,b)$.
Furthermore, if the method is an oblivious one-step gradient method on $g$, i.e.,
@@
Inductive step. Assume (4) holds for all indices up to $t$. By the span property (2), there exist scalars $\{\beta_{j}^{(t)}\}_{j=0}^{t}$ such that
@@
(If the method is oblivious span, we take $\{\beta_{j}^{(t)}\}$ to be exactly the pre-scheduled coefficients specified by the method.) Subtracting $x^{\star}$ and using (3) gives
@@
Obliviousness and instance-independence. If the method is an oblivious span method (Definition 3.11), then the coefficients $\{\beta_{j}^{(t)}\}_{j=0}^{t}$ in (5) are fixed in advance (they may depend on $t$ and on known problem parameters, but not on the realized oracle replies). Hence the recursion that constructs $p_{t+1}$ from $(p_{0},\ldots,p_{t})$ uses only these fixed coefficients. Therefore, for each $t$, the resulting polynomial $p_{t}$ is determined solely by the method’s fixed coefficients (and by $t$), and in particular does not depend on $(H,b)$.
@@ 4 Oracle Complexity Lower Bounds
4.1 A Non-Degenerate Lifting Construction
A common concern with oracle lower bounds is that the hard instances may be degenerate from a multiobjective perspective—for example, the objectives may coincide up to constants, or the Pareto set may collapse to a singleton—so that the lower bound reflects a scalar artifact rather than a genuinely multiobjective phenomenon. To rule this out, we introduce a simple but robust lifting device: it embeds an arbitrary scalar function $g$ defined on a subspace $V$ into an $m$-objective instance on $\mathbb{R}^{d}=V\oplus W$ by adding objective-dependent quadratic terms on a complementary subspace $W$. The resulting objectives are distinct, the Pareto set is non-singleton, and the Pareto stationarity gap $\mathcal{G}$ admits an explicit decomposition showing that it dominates the scalar stationarity measure $\|\nabla g\|$. This “gap domination” property is the mechanism that transfers first-order lower bounds from the scalar problem to the multiobjective setting (including nonconvex $g$, since convexity is not used in the domination argument).
Theorem 4.1 (Non-degenerate lifting and stationarity-gap domination).
Fix integers $m\geq 2$ and $d\geq 1$. Let $\mathbb{R}^{d}=V\oplus W$ be an orthogonal direct sum with $\dim(W)\geq m-1$, and pick $m$ affinely independent points $a_{1},\dots,a_{m}\in W$. Let $g:V\to\mathbb{R}$ be $C^{1}$ with $L_{g}$-Lipschitz gradient, and fix $\gamma>0$. Define for $i=1,\dots,m$:
$f_{i}(x):=g(x_{V})+\frac{\gamma}{2}\|x_{W}-a_{i}\|^{2},\qquad x=x_{V}+x_{W}.$ (6)
Then:
  1. (i)
    The objectives $\{f_{i}\}_{i=1}^{m}$ are distinct.
  2. (ii)
    Each $f_{i}$ is $L$-smooth with $L=\max\{L_{g},\gamma\}$. If $g$ is convex (resp. $\mu$-strongly convex) and $\gamma\geq 0$ (resp. $\gamma\geq\mu$), then each $f_{i}$ is convex (resp. $\mu$-strongly convex).
  3. (iii)
    For all $x\in\mathbb{R}^{d}$,
    $\mathcal{G}(x)^{2}=\|\nabla g(x_{V})\|^{2}+\gamma^{2}\,\mathrm{dist}\!\left(x_{W},\mathrm{conv}\{a_{1},\dots,a_{m}\}\right)^{2},$ (7)
    in particular $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$.
  4. (iv)
    If $g$ is convex and admits a minimizer (i.e., $\arg\min g\neq\emptyset$), then the (weak) Pareto set equals
    $\mathcal{P}=(\arg\min g)\times\mathrm{conv}\{a_{1},\dots,a_{m}\}.$
    In particular, $\mathrm{conv}\{a_{1},\dots,a_{m}\}$ is a non-singleton polytope. If $g$ is $\mu$-strongly convex, then $\arg\min g=\{x_{V}^{\star}\}$ is a singleton and $\mathcal{P}=\{x_{V}^{\star}\}\times\mathrm{conv}\{a_{1},\dots,a_{m}\}$.
  5. (v)
    Regardless of convexity of $g$, the Pareto-stationary set of the lifted instance satisfies
    $\mathcal{S}:=\{x:\mathcal{G}(x)=0\}=\{x_{V}\in V:\nabla g(x_{V})=0\}\times\mathrm{conv}\{a_{1},\dots,a_{m}\}.$
Consequently, any oracle lower bound on $\|\nabla g(u^{(t)})\|$ for a scalar first-order method applied to $g$ transfers verbatim to a lower bound on $\mathcal{G}(x^{(t)})$ for the lifted MOO instance.
@@ 4.1 A Non-Degenerate Lifting Construction
Throughout, we use the orthogonal decomposition $\mathbb{R}^{d}=V\oplus W$ and write $x=x_{V}+x_{W}$ with $x_{V}\in V$ and $x_{W}\in W$. Differentiating (6) yields, for each $i$,
$\nabla f_{i}(x)=\bigl(\nabla g(x_{V}),\ \gamma(x_{W}-a_{i})\bigr).$ (8)
(i) Distinctness of objectives. Because the points $a_{1},\ldots,a_{m}$ are affinely independent, in particular they are pairwise distinct. Fix $i\neq j$ and take any $x_{V}\in V$ with $x_{W}=a_{i}$. Then
$f_{i}(x_{V}+a_{i})=g(x_{V}),\qquad f_{j}(x_{V}+a_{i})=g(x_{V})+\frac{\gamma}{2}\|a_{i}-a_{j}\|^{2}>g(x_{V}),$
so $f_{i}\not\equiv f_{j}$. Hence the objectives are distinct.
(ii) Smoothness and (strong) convexity. We first verify smoothness. Since $g$ has $L_{g}$-Lipschitz gradient on $V$, for all $u,v\in V$, $\|\nabla g(u)-\nabla g(v)\|\leq L_{g}\|u-v\|$. Using (8), we obtain for any $x,y\in\mathbb{R}^{d}$,
$\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2}$ $=\|\nabla g(x_{V})-\nabla g(y_{V})\|^{2}+\gamma^{2}\|x_{W}-y_{W}\|^{2}$
$\leq L_{g}^{2}\|x_{V}-y_{V}\|^{2}+\gamma^{2}\|x_{W}-y_{W}\|^{2}$
$\leq\max\{L_{g}^{2},\gamma^{2}\}\bigl(\|x_{V}-y_{V}\|^{2}+\|x_{W}-y_{W}\|^{2}\bigr)$
$=\max\{L_{g}^{2},\gamma^{2}\}\,\|x-y\|^{2},$
where we used orthogonality of $V$ and $W$ in the last step. Taking square roots yields $\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq\max\{L_{g},\gamma\}\|x-y\|$, hence each $f_{i}$ is $L$-smooth with $L=\max\{L_{g},\gamma\}$.
If $g$ is convex, then $x\mapsto g(x_{V})$ is convex on $\mathbb{R}^{d}$ (as a composition of the linear projection $x\mapsto x_{V}$ with a convex function), and $x\mapsto\frac{\gamma}{2}\|x_{W}-a_{i}\|^{2}$ is convex for $\gamma\geq 0$; therefore $f_{i}$ is convex. If $g$ is $\mu$-strongly convex on $V$, then $x\mapsto g(x_{V})$ is $\mu$-strongly convex along directions in $V$, while $x\mapsto\frac{\gamma}{2}\|x_{W}-a_{i}\|^{2}$ is $\gamma$-strongly convex along directions in $W$; consequently, for all $x,y\in\mathbb{R}^{d}$,
$f_{i}(y)\geq f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{\min\{\mu,\gamma\}}{2}\|y-x\|^{2},$
so $f_{i}$ is $\min\{\mu,\gamma\}$-strongly convex. In particular, if $\gamma\geq\mu$ then each $f_{i}$ is $\mu$-strongly convex.
(iii) Stationarity-gap formula and domination. For any $\lambda\in\Delta^{m}$, summing (8) gives
$\sum_{i=1}^{m}\lambda_{i}\nabla f_{i}(x)=\left(\nabla g(x_{V}),\ \gamma\Bigl(x_{W}-\sum_{i=1}^{m}\lambda_{i}a_{i}\Bigr)\right).$
By orthogonality of $V$ and $W$,
$\left\|\sum_{i=1}^{m}\lambda_{i}\nabla f_{i}(x)\right\|^{2}=\|\nabla g(x_{V})\|^{2}+\gamma^{2}\left\|x_{W}-\sum_{i=1}^{m}\lambda_{i}a_{i}\right\|^{2}.$
The first term is independent of $\lambda$, and $\{\sum_{i}\lambda_{i}a_{i}:\lambda\in\Delta^{m}\}=\mathrm{conv}\{a_{1},\ldots,a_{m}\}$, hence
$\mathcal{G}(x)^{2}$ $=\min_{\lambda\in\Delta^{m}}\left\|\sum_{i=1}^{m}\lambda_{i}\nabla f_{i}(x)\right\|^{2}$
$=\|\nabla g(x_{V})\|^{2}+\gamma^{2}\,\mathrm{dist}\!\left(x_{W},\mathrm{conv}\{a_{1},\ldots,a_{m}\}\right)^{2},$
which is (7). In particular, $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$.
(iv) Pareto set in the convex case. Assume $g$ is convex and admits a minimizer $x_{V}^{\star}\in\arg\min g$. By Proposition 3.8, weak Pareto optima coincide with minimizers of some scalarization $f_{\lambda}=\sum_{i=1}^{m}\lambda_{i}f_{i}$. For any $\lambda\in\Delta^{m}$,
$f_{\lambda}(x_{V},x_{W})$ $=g(x_{V})+\frac{\gamma}{2}\sum_{i=1}^{m}\lambda_{i}\|x_{W}-a_{i}\|^{2}$
$=g(x_{V})+\frac{\gamma}{2}\left\|x_{W}-\bar{a}(\lambda)\right\|^{2}+\mathrm{const}(\lambda),\qquad\bar{a}(\lambda):=\sum_{i=1}^{m}\lambda_{i}a_{i},$
where $\mathrm{const}(\lambda)$ is independent of $(x_{V},x_{W})$. Thus minimization decouples: the $V$-component is minimized by any $x_{V}\in\arg\min g$, and the quadratic in $x_{W}$ is uniquely minimized at $x_{W}=\bar{a}(\lambda)$. Hence
$\arg\min_{x\in\mathbb{R}^{d}}f_{\lambda}(x)=(\arg\min g)\times\{\bar{a}(\lambda)\}.$
As $\lambda$ ranges over $\Delta^{m}$, $\bar{a}(\lambda)$ ranges over $\mathrm{conv}\{a_{1},\ldots,a_{m}\}$, yielding
$\mathcal{P}=(\arg\min g)\times\mathrm{conv}\{a_{1},\dots,a_{m}\}.$
Since $m\geq 2$ and the $a_{i}$ are affinely independent, $\mathrm{conv}\{a_{i}\}$ is a non-singleton polytope. If $g$ is $\mu$-strongly convex, then $\arg\min g=\{x_{V}^{\star}\}$ is a singleton and the same description reduces to $\mathcal{P}=\{x_{V}^{\star}\}\times\mathrm{conv}\{a_{1},\dots,a_{m}\}$.
(v) Pareto-stationary set for general $g$. By (7), $\mathcal{G}(x)=0$ holds if and only if $\nabla g(x_{V})=0$ and $\mathrm{dist}(x_{W},\mathrm{conv}\{a_{i}\})=0$, i.e., $x_{W}\in\mathrm{conv}\{a_{1},\ldots,a_{m}\}$. Therefore,
$\mathcal{S}=\{x:\mathcal{G}(x)=0\}=\{x_{V}\in V:\nabla g(x_{V})=0\}\times\mathrm{conv}\{a_{1},\ldots,a_{m}\}.$
Transfer claim. The pointwise domination $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ from (iii) implies that for any iterate sequence $(x^{(t)})_{t\geq 0}$,
$\mathcal{G}(x^{(t)})\geq\|\nabla g(x_{V}^{(t)})\|\qquad\text{for all }t.$
Hence any lower bound on $\|\nabla g(\cdot)\|$ along the projected iterates transfers immediately to a lower bound on $\mathcal{G}(\cdot)$ along the lifted iterates. ∎
4.2 Scalar hard instances: polynomial-degree and geometric hardness
We will use two types of scalar lower bounds: (i) quadratic lower bounds for span/oblivious methods that reduce to extremal polynomial problems; and (ii) geometric (zero-chain) lower bounds for finding stationary points under a bounded initial value gap, due to [6, 7].
4.2.1 Quadratic hard instances for span/oblivious methods
We first construct scalar quadratic hard instances tailored to span/oblivious first-order methods. For quadratics, these methods generate iterates that admit a polynomial/Krylov representation, which enables explicit finite-horizon lower bounds on $\min_{t\leq T}\|\nabla g(x^{(t)})\|$ and separates the behavior of the different method classes. Lemma 4.2 below states the specific scalar quadratic instances and bounds that will serve as base ingredients for our multiobjective lower bounds via the lifting theorem.
Lemma 4.2 (Hard scalar quadratic instances).
Fix $T\in\mathbb{N}$ and $R>0$.
  1. (a)
    (Strongly convex case) Fix $L\geq\mu>0$ and $\kappa:=L/\mu$. For any span first-order method making $T$ oracle calls on quadratics, there exist a dimension $d\geq T+1$, an $L$-smooth $\mu$-strongly convex quadratic $g$ and an initialization with $\|x^{(0)}-x^{\star}\|=R$ such that the $T$-th iterate satisfies
    $\|\nabla g(x^{(T)})\|\;\geq\;\mu\,R\,\frac{2}{\rho^{T}+\rho^{-T}}\;\geq\;\mu\,R\left(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\right)^{T},\qquad\rho:=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$
  2. (b)
    (Convex, oblivious one-step) Fix $L>0$. For any oblivious one-step gradient method run for $T$ steps with step sizes $\alpha_{t}\in[0,1/L]$, there exists an $L$-smooth convex quadratic $g$ and an initialization with $\|x^{(0)}-x^{\star}\|=R$, where $x^{\star}\in\arg\min g$ is chosen as a closest minimizer to $x^{(0)}$, such that
    $\min_{0\leq t\leq T}\|\nabla g(x^{(t)})\|\;\geq\;\frac{LR}{4(T+1)}.$
  3. (c)
    (Convex, oblivious span: polynomial-degree) Fix $L>0$. For any oblivious span first-order method run for $T$ steps, there exists an $L$-smooth convex quadratic $g$ and an initialization with $\|x^{(0)}-x^{\star}\|=R$, where $x^{\star}\in\arg\min g$ is chosen as a closest minimizer to $x^{(0)}$, such that
    $\|\nabla g(x^{(T)})\|\;\geq\;\frac{LR}{2(T+1)^{2}}.$
Proof.
We prove each part separately.
@@ 4.2.1 Quadratic hard instances for span/oblivious methods
(a) Strongly convex case. Let $g(x)=\frac{1}{2}x^{\top}Hx-b^{\top}x$ with $H\succ 0$. For any iterate sequence $\{x^{(t)}\}$ produced by a span first-order method on a quadratic, the error admits the polynomial representation (Lemma 3.15):
@@
where the last inequality follows from $\frac{2}{a+a^{-1}}\geq a^{-1}$ for $a\geq 1$.
(b) Convex oblivious one-step case. This is precisely Corollary A.1.4 in Appendix A.1 (with $t=T$).
(c) Convex oblivious span polynomial-degree case. Fix an arbitrary oblivious span method and consider any $L$-smooth convex quadratic $g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x$ with $\sigma(H)\subset[0,L]$. By Lemma 3.15 (oblivious case), there exists a polynomial $p_{T}$ with $\deg p_{T}\leq T$ and $p_{T}(0)=1$ that depends only on the method (and on $T$), such that
$x^{(T)}-x^{\star}=p_{T}(H)\,(x^{(0)}-x^{\star}),$
and hence
$\nabla g(x^{(T)})=H\,p_{T}(H)\,(x^{(0)}-x^{\star}).$
Let $\|x^{(0)}-x^{\star}\|=R$. For any fixed $\zeta\in[0,L]$, choose $H$ diagonal with a single eigenvalue $\zeta$ (and all remaining eigenvalues equal to 0) and align $x^{(0)}-x^{\star}$ with the corresponding eigenvector. Then
$\|\nabla g(x^{(T)})\|\;=\;R\,|\zeta\,p_{T}(\zeta)|.$
Since $p_{T}$ is fixed independently of $\zeta$ (obliviousness), taking the worst case over $\zeta\in[0,L]$ yields
$\sup_{\begin{subarray}{c}g\ \text{$L$-smooth convex quadratic}\\ \|x^{(0)}-x^{\star}\|=R\end{subarray}}\|\nabla g(x^{(T)})\|\;\geq\;R\,\max_{\zeta\in[0,L]}|\zeta p_{T}(\zeta)|.$
Finally, taking the infimum over oblivious span methods is lower bounded by the infimum over all admissible polynomials, and Appendix A.2 gives
This yields $\|\nabla g(x^{(T)})\|\geq\frac{LR}{2(T+1)^{2}}$ as claimed. ∎
4.2.2 Geometric stationary-point lower bounds under value-gap bounds
We next state the geometric scalar stationary-point lower bounds that drive our new convex and nonconvex MOO results.
Lemma 4.3 (Convex $L$-smooth stationary-point lower bound under bounded value gap [7, Theorem 1]).
Fix the initial point $x^{(0)}=0$. A deterministic first-order method $\mathcal{A}$ is any procedure that, for $t=0,1,2,\dots$, adaptively chooses query points
$x^{(t+1)}=\Phi_{t}\!\left(\{(g(x^{(s)}),\nabla g(x^{(s)}))\}_{s=0}^{t}\right),$
for some deterministic mappings $\Phi_{t}$, where the oracle returns $(g(x),\nabla g(x))$ at each query.
Fix $L>0$ and $\Delta>0$. For any deterministic first-order method $\mathcal{A}$ and any $T\in\mathbb{N}$, there exist a dimension $d\geq 2T$ and a convex function $g:\mathbb{R}^{d}\to\mathbb{R}$ with $L$-Lipschitz gradient such that
$g(x^{(0)})-\inf_{x\in\mathbb{R}^{d}}g(x)\leq\Delta,$
and the iterates $\{x^{(t)}\}_{t=0}^{T}$ generated by $\mathcal{A}$ on $g$ satisfy
$\min_{0\leq t\leq T}\|\nabla g(x^{(t)})\|\;\geq\;\frac{\sqrt{L\Delta}}{4(T+1)}.$
Equivalently, any deterministic first-order method requires $T=\Omega\!\left(\sqrt{L\Delta}/\varepsilon\right)$ oracle calls in the worst case to guarantee an output $x$ with $\|\nabla g(x)\|\leq\varepsilon$ over this function class.
Proof.
Carmon, Duchi, Hinder, and Sidford [7, Theorem 1] prove the following: for any $\varepsilon>0$ and any deterministic first-order method $\mathcal{A}$, there exists a convex function $f$ with $L$-Lipschitz gradient and $f(x^{(0)})-\inf_{x}f(x)\leq\Delta$ such that $\mathcal{A}$ must make at least $\frac{\sqrt{L\Delta}}{4\varepsilon}$ first-order oracle calls to guarantee finding a point $x$ with $\|\nabla f(x)\|\leq\varepsilon$. (Their statement is given for an orthogonally invariant function class $K_{1}(\Delta,L)$, and the reduction from zero-respecting to general deterministic methods is handled via a resisting-oracle / rotation argument, which increases the dimension to at most $2T$.)
Now fix $T\in\mathbb{N}$ and set $\varepsilon:=\frac{\sqrt{L\Delta}}{4(T+1)}$. Then $\frac{\sqrt{L\Delta}}{4\varepsilon}=T+1$, so any method using only $T$ oracle calls cannot guarantee $\|\nabla g(x)\|\leq\varepsilon$ on the above convex $L$-smooth class. Therefore, for the corresponding worst-case instance $g$, all iterates up to time $T$ satisfy $\|\nabla g(x^{(t)})\|>\varepsilon$, and hence
$\min_{0\leq t\leq T}\|\nabla g(x^{(t)})\|\geq\varepsilon=\frac{\sqrt{L\Delta}}{4(T+1)}.$
Rearranging yields the equivalent complexity form $T=\Omega(\sqrt{L\Delta}/\varepsilon)$. ∎
Lemma 4.4 (Nonconvex $L$-smooth stationary-point lower bound (Lipschitz gradient only) [6, Theorem 1]).
Fix the initial point $x^{(0)}=0$. A deterministic first-order method $\mathcal{A}$ is any procedure that, for $t=0,1,2,\dots$, adaptively chooses query points
$x^{(t+1)}=\Phi_{t}\!\left(\{(g(x^{(s)}),\nabla g(x^{(s)}))\}_{s=0}^{t}\right),$
for some deterministic mappings $\Phi_{t}$, where the oracle returns the pair $(g(x),\nabla g(x))$ at each query.
There exists a universal constant $c>0$ such that for any $L>0$, $\Delta>0$, any deterministic first-order method $\mathcal{A}$, and any $T\in\mathbb{N}$, there exist a dimension $d\geq 2T+1$ and an $L$-smooth (not necessarily convex) function $g:\mathbb{R}^{d}\to\mathbb{R}$ satisfying
$g(x^{(0)})-\inf_{x\in\mathbb{R}^{d}}g(x)\leq\Delta,$
such that the iterates $\{x^{(t)}\}_{t=0}^{T}$ generated by $\mathcal{A}$ on $g$ obey
$\min_{0\leq t\leq T}\|\nabla g(x^{(t)})\|\;\geq\;c\,\sqrt{\frac{L\Delta}{T+1}}.$
Equivalently, any deterministic first-order method requires $T=\Omega(L\Delta/\varepsilon^{2})$ oracle calls in the worst case to guarantee an output $x$ with $\|\nabla g(x)\|\leq\varepsilon$.
Proof.
Carmon, Duchi, Hinder, and Sidford prove (see [6, Theorem 1] specialized to $p=1$) that there exists a universal constant $C>0$ such that for every $\varepsilon>0$, every deterministic first-order method $\mathcal{A}$ requires at least
$T_{\varepsilon}\;\geq\;C\,\frac{L\Delta}{\varepsilon^{2}}$
oracle calls in the worst case over the class of $L$-smooth functions with initial gap at most $\Delta$, i.e., there exists an $L$-smooth (possibly nonconvex) $g$ with $g(x^{(0)})-\inf g\leq\Delta$ such that $\|\nabla g(x^{(t)})\|>\varepsilon$ for all $t<T_{\varepsilon}$.
Now fix $T\in\mathbb{N}$ and set $\varepsilon:=\sqrt{\frac{CL\Delta}{T+1}}$. Then $T+1<C\frac{L\Delta}{\varepsilon^{2}}$, hence $T<T_{\varepsilon}$ for the corresponding worst-case instance $g$, which implies $\|\nabla g(x^{(t)})\|>\varepsilon$ for all $t\leq T$ and therefore
$\min_{0\leq t\leq T}\|\nabla g(x^{(t)})\|\;\geq\;\varepsilon\;=\;\sqrt{\frac{CL\Delta}{T+1}}.$
Renaming $c:=\sqrt{C}$ yields the stated bound. The equivalence to $T=\Omega(L\Delta/\varepsilon^{2})$ follows by rearranging the inequality. ∎
4.3 Strongly Convex MOO: A Tight Linear Rate
We now state the corresponding multiobjective lower bound in the strongly convex regime.
Theorem 4.5 (Strongly Convex Lower Bound).
Fix $L\geq\mu>0$, $T\in\mathbb{N}$, $\kappa:=L/\mu$, and $R>0$. Consider any span first-order method that makes $T$ first-order oracle calls when applied to a quadratic instance. Then there exist integers $m\geq 2$ and $d\geq(T+1)+(m-1)$, an MOO instance $(f_{1},\dots,f_{m}):\mathbb{R}^{d}\to\mathbb{R}^{m}$ with $m$ objectives each $L$-smooth and $\mu$-strongly convex, and an initialization $x^{(0)}$ with $\mathrm{dist}(x^{(0)},\mathcal{P})=R$ such that the $T$-th iterate produced by the method satisfies
$\mathcal{G}\bigl(x^{(T)}\bigr)$ $\;\geq\;\mu\,R\;\frac{2}{\rho^{T}+\rho^{-T}}$
$\;\geq\;\mu\,R\left(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\right)^{T},\qquad\rho:=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$
In particular, any such method has worst-case iteration complexity
$\Omega\!\bigl(\sqrt{\kappa}\log(1/\varepsilon)\bigr)$
to reduce the Pareto stationarity gap below $\varepsilon$.
Proof.
By Lemma 4.2(a), there exists a dimension $\dim(V)\geq T+1$ and an $L$-smooth $\mu$-strongly convex quadratic $g$ on a subspace $V$ that forces the stated lower bound on $\|\nabla g(u^{(T)})\|$ for any span method (with $\|u^{(0)}-u^{\star}\|=R$). Apply the lifting construction of Theorem 4.1 with $\gamma=\mu$ to obtain an MOO instance $(f_{1},\dots,f_{m})$ with $\mathrm{dist}(x^{(0)},\mathcal{P})=R$ and
$\mathcal{G}(x^{(T)})\;\geq\;\|\nabla g(x_{V}^{(T)})\|.$
Since the $V$-component of the iterate sequence produced by a span method on the lifted instance satisfies the span property with respect to $g$, the scalar lower bound transfers, proving the claim. The iteration-complexity statement follows by solving $\mu R\bigl((\sqrt{\kappa}-1)/(\sqrt{\kappa}+1)\bigr)^{T}\leq\varepsilon$ for $T$. ∎
4.4 Convex MOO: Oblivious vs. Adaptive Methods